OpenStudy (anonymous):

square root of 6 times square root of 216

OpenStudy (anonymous):

36

OpenStudy (amistre64):

sqrt(6^1) * sqrt(6^3) = sqrt(6^1 * 6^3) sqrt(6^4) = 6^2 = 36.... yep